# Dynamically Augmented CVaR for MDPs

### v3

@@ Abstract
This paper studies optimization of Conditional Value-at-Risk (CVaR) for Markov Decision Processes (MDPs) with finite state and action sets. It introduces the Dynamically augmented CVaR (DCVaR) risk measure and provides an algorithm for its optimization. This paper investigates a specially defined Robust MDP (RMDP), in which the state space is augmented with the tail risk level. This RMDP, which we call the Dynamically augmented RMDP (DRMDP), was introduced to the literature for calculations of optimal CVaR values by value iteration more than ten years ago, but, as was understood later, these value iterations compute lower bounds of minimal static CVaRs. DCVaR is defined as a time consistent version of the static CVaR, and it is a lower bound of the static CVaR. It also can be considered as a dynamic version of the nested CVaR. This paper provides an algorithm constructing a policy optimizing DCVaR of total discounted costs. The correctness of this algorithm is proved by studying a special mass transfer problem. The results on RMDPs needed for this paper are provided in the appendix.
@@ 1 Introduction
Conditional Value-at-Risk (CVaR), also sometimes called Average Value-at-Risk (AVaR), is one of the most important and popular risk measures. For dynamic problems, it is possible to consider the so-called static CVaR, when the CVaR value is defined for each policy, and the goal is to find a policy optimizing this value. However, there are problems with this approach dealing with computational complexity of finding optimal policies and time inconsistency of static CVaRs. The issue of time consistency of risk measures is discussed in several works, including [21, 30, 31, 33, 34, 35]. An alternative approach is to consider the nested CVaR, which is defined via a Robust MDP (RMDP) whose state space coincides with the original state space of the problem. Another approach, which is also often used in practice, is to consider an RMDP, whose states are pairs consisting of the original state augmented with the risk level. In other words, the state is dynamically augmented by a number between 0 and 1 representing the tail risk level.
@@
The first group of the results in this paper establishes the relation between the static CVaR values and the DRMDP. First we observe that there is an optimal nonrandomized policy minimizing static CVaR, and, for a nonrandomized policy for an MDP, its static CVaR is equal to the worst possible expected outcome if the Decision Maker (DM) plays this policy in the DRMDP. This observation clarifies the gap described in Hau et al. [15] and illustrates time inconsistency of the static CVaR for MDPs because, in order to implement this worst possible outcome, the second player in the DRMDP, whom we call Nature, may have to play a policy, which knows future decisions of the DM. A more natural assumption is that Nature plays its optimal policy, and this assumption leads to the definition of the DCVaR provided in this paper. This definition implies that value iterations proposed in Chow et al. [4] converge to minimal values of DCVaR.
@@
Classic methods for Markov Decision Processes (MDPs) deal with risk-neutral objectives. In addition, there is a significant literature on risk-sensitive optimization initiated by Howard and Matheson [18] and dealing mostly with optimization of expectations of utility functions rather than with optimization of expected total costs or average costs per unit time. Pioneering work by Markowitz [22] on mean-variance optimization for portfolio selection influenced research on MDPs with mean-variance criteria [11, 37], but variance of the total costs is not a convenient characteristic to deal with for sequential problems, and it is not a good risk measure, because variance is affected by losses and gains in the same way. CVaR is an attractive alternative to variance both because of its practical importance and rich mathematical properties; see [26, 27, 28, 36]. CVaR is a risk measure that has gained popularity in various applications including finance. It has the ability to safeguard a DM from the tail events by focusing on an average of the largest losses and by doing so provides a more comprehensive view for risk management than threshold-based risk measures such as Value-at-Risk(VaR). CVaR is also known as expected shortfall in financial literature and has great importance in terms of financial regulation.
@@
We summarize the organization of this paper and its main results. Preliminary information on CVaR and MDPs is provided in Section 2. In Section 3 we prove the existence of a nonrandomized policy minimizing CVaR of the total discounted costs. Then we define the DRMDP, which is an RMDP in which the DM chooses actions, and Nature assigns risk levels. Policies for the original MDP are particular policies of the DM in the DRMDP when the DM does not observe past and current tail risk levels, except for the initial one, and we call these policies risk-independent. We show that for a nonrandomized risk-independent policy played by the DM, CVaR of total discounted costs is equal to the worst expected total discounted costs the DM can get by playing this policy in the DRMDP. This result illustrates time inconsistency of the static CVaR and implies that the minimal CVaR value for the MDP is equal to the minimal value for the class of nonrandomized risk-independent policies in the DRMDP. This equality illustrates a positive gap between the optimal static CVaR value and the limit of value iterations, which was discovered in Hau et al. [15]. Section 4 defines DCVaR, describes its properties, and introduces a DRMDP1, which is obtained from the DRMDP by modifying one-step costs and transition probabilities. The payoff and value functions for the DRMDP1 are equal to these functions for the DRMDP multiplied by tail risk levels. The sets of optimal policies for the DRMDP and DRMP1 coincide if the initial tail risk level is positive. The advantage of the DRMDP1 is that its value function is concave in the tail risk level.
@@
Section 5 introduces the algorithm for constructing a nonrandomized optimal policy minimizing DCVaR, which was announced in [5]. Sections 6–8 deal with proving that this algorithm constructs an optimal policy. Section 6 describes the properties of the mass transfer problem relevant to optimal solutions for Nature. Section 7 describes the properties of optimal values and decisions for Nature. Section 8 proves correctness of the algorithm. Section 9 provides extensions to random cost. Appendix A provides results on RMDPs used in this paper.
@@ 2 Preliminaries: CVaR and MDPs
where is the distribution function of In addition, if this expectation exists, and where by definition; see [26, 27, 28] and [36, Chapter 6] for additional details regarding CVaR.
@@
where , and is the set of random variables on the probability space defined as and is called the dual CVaR risk envelope. In this paper, we sometimes write a constant instead of functions equal to this constant, and comparisons between random variables are usually understood -a.s.
Let be a sub -algebra of that is, is a -algebra on and A -measurable random variable is called a random tail risk level. Then the conditional CVaR given the -algebra and a random tail risk level is defined [25, Definition 17] as the -measurable random variable
where is the set of nonnegative -measurable random variables on such that and According to the CVaR decomposition theorem by Pflug and Pichler ([25], Lemma 22),
@@
An MDP is a model of a natural family of controlled stochastic sequences. It is a tuple where is a set of states, is a set of actions, and for each state there is a nonempty set of available actions is a bounded cost function, and is a transition probability. The time is discrete, and, if at some time instance an action is selected at a state , then the system moves to the next state according to and the cost is collected, allowing the cost to depend on the next state. In this paper we always assume that and are finite sets.
@@
The risk-neutral approach is to consider the expected total discounted cost For risk level and a finite horizon or an infinite horizon let us consider the objective function where , and is the random variable defined on the probability space in formulae (2.6) or (2.7). The value is sometimes called a static CVaR. Computing optimal policies minimizing the static CVaR is a hard problem [2, Remark 5.1].
Chow et al. [4] changed the original problem formulation with the state space to the formulation when the states of the problem consist of state-risk pairs where is the original state and is the tail risk level. In other words, the state space is augmented in [4] with the tail risk level, and the new state space is Motivated by the Pflug-Pichler CVaR decomposition theorem [25], Chow et al. [4] introduced the Robust MDP (RMDP) with the state space which is formally defined in Section 3 and called the Dynamically augmented RMDP (DRMDP). Hau et al. showed that the value of DRMDP is a lower bound of the minimal CVaR values for the original MDP. In Section 4 we introduce the Dynamically augmented CVaR (DCVaR) objective function, which is a lower bound of CVaR, and the value of DRMDP is equal to the minimal value of DCVaR.
In the conclusion of this section, we recall two concepts for MDPs relevant to more general objective criteria than expected total costs: MDPs with general objective functions and abstract dynamic programming. According to [7], a general objective function is a function such that, if for and then This definition is consistent with the notion of a law-invariant risk measure [25, 36]. The value function is A policy is optimal for the initial state if A policy is optimal if it is optimal for all initial states
@@ 3 Static CVaR and Robust MDPs
The main results of this section are: (i) the existence of nonrandomized policies minimizing static CVaR (Theorem 3.1), and (ii) the minimal value of static CVaR is the optimal minimal guaranteed expected total discounted payoff for the DM in the RMDP introduced in Chow et al. [4], if the DM plays policies from the MDP; see Corollary 3.4 below. This corollary provides a game-theoretic interpretation of the gap between the optimal value of CVaR and optimal value of this RMDP discovered in Hau et al. [15].
@@
be the optimal value of the static CVaR. The following theorem states the existence of nonrandomized optimal policies minimizing static CVaRs and convergence of finite-horizon optimal values of static CVaRs to the optimal infinite-horizon value. We recall that is fixed. We also recall that throughout the entire paper if and if
@@
For every or for there exist a nonrandomized optimal policy for the CVaR optimization problem for which for all In addition,
@@
We observe that it is sufficient to proof that for every there exists a nonrandomized optimal policy for the initial state Indeed, let be the set of all nonrandomized -horizon policies. If is optimal for an initial state then the nonrandomized policy is optimal, where Let us fix and prove the existence of Let us start with In this case the set is finite. Therefore, the strategic measure for an arbitrary policy can be presented as a convex combination of strategic measures for nonrandomized policies [6, Theorem 1]: there are numbers such that and Since taking CVaR of a mixture of distributions is a concave operation [24, Proposition 3.2],
@@
Let us consider the infinite-horizon problem. Also, in the rest of this proof, let us set for all If we add a constant to the cost function then the all CVaR values for an -horizon problem will be changed by the constant Thus, this addition preserves sets of optimal policies. Since the cost function is bounded, without loss of generality we can assume that the function is nonnegative. Let us assume that and let is an upper bound of the cost function, that is, for all and
@@
Let us construct a nonrandomized optimal infinite-horizon policy Let be nonrandomized optimal -horizon policies. Since the sets and where are finite, it is possible to choose a sequence of integers such that for all We define for all
Now let for some there is an increasing sequence of natural numbers such that for each history policies make the same decisions We select a subsequence of the sequence such that for each history all policies make the same decision at the step By repeating this procedure and taking we construct a nonrandomized policy
@@
Let us consider the RMDP introduced in Chow et al. [4], which we call the Dynamically augmented RMDP (DRMDP). The DRMDP is defined by a tuple where the state space is action space is uncertainty space is with being the number of states in action sets for the DM at states are uncertainty sets for Nature are
where, for
one-step costs for and transition probabilities
where is the Dirac measure on the interval concentrated at the point A DRMDP is a particular case of the RMDP model described in Appendix A, when the sets set-valued mapping one-step cost function and transition probability have specific forms. In particular, states for DRMDPs are denoted by The sets of policies for the DM and Nature are denoted by and respectively. We shall use notation where and for the expected total discounted payoffs over the horizon with the terminal payoff which is continuous in As explained after formulae (A.2) and (A.4), the optimal values are continuous functions. Since is a finite set, this means that the functions are continuous in
We observe that that is, the set of all policies for the original MDP is the subset of the set of policies for the DM in the DRMDP. We call policies from risk-independent. The fundamental difference between a policy and a policy is that the policy does not have information about the current and past tail risk levels except the initial tail risk level which we consider to be fixed.
The following theorem states that for each risk-independent policy for the DM, Nature can play a policy maximizing DM’s losses. The proofs of Theorems 3.2 and A.2 are based on the well-known fact that, if in a two-player stochastic game a policy of one player is fixed, then another player deals with an MDP. In general, the states of this MDP are histories in the game. However, the state space of this MDP can be simplified, if the fixed policy belongs to a special class. For example, if the fixed policy is stationary, then another player deals with an MDP whose states are the states of the stochastic game. There are differences between the proofs of Theorems 3.2 and A.2. The proof of Theorem A.2 follows from the validity of sufficient conditions for the existence of optimal policies for MDPs with setwise continuous transition probabilities [32, 9], which takes place because all the sets are finite. The proof of Theorem 3.2 follows from the validity of sufficient conditions for the existence of optimal policies for MDPs with weakly continuous transition probabilities [32, 8], which takes place because the set and all the sets are finite and because of weak continuity of the transition probability The remaining statements of this section deal with the DRMDP.
@@
For each risk-independent policy for the DM and for each or there exists a nonrandomized policy which depends on such that
@@
If the DM plays a risk-independent policy Nature deals with an MDP with states The set of available actions at each state is where and The one-step reward for Nature is and this function is bounded and continuous, where If an action is chosen at state then the next state is with the probability if the following conditions hold: (a) (b) (c) (d) Other transitions are impossible since The transition probability is weakly continuous in because is a bounded continuous function and and are distributions on finite sets depending on the finite numbers of conditions. We consider the expected total discounted rewards for this problem with the discount factor The initial state is where which is the initial tail risk level, and This MDP satisfies the weakly continuous conditions, which implies the existence of optimal Markov policies for finite-horizon problems and optimal deterministic policies for infinite-horizon problems [32] or [9, Theorem 2]. These optimal policies define the policy for Nature whose existence is stated in the theorem. ∎
@@
Let us consider the MDP defined at the beginning on the proof of Theorem 3.2 for We fix and compute an optimal policy for Nature. If at time Nature is at state of this MDP, then the optimal 1-step value for Nature for is
there is an optimal 1-step Markov policy [9, Theorem 2] and, in view of Berge’s maximum theorem, the function is continuous. This means that this function is continuous in in since all other variables in take a finite number of values. Nature chooses the optimal action without using any knowledge of the previous states and actions. This is true for because the optimal value of does not depend on the distributions of future actions and states following the state
Let us make the induction assumption that for some integer there is a Markov policy such that this policy is optimal for the horizon and the value function is continuous, which means that it is continuous in where for Then
where: (i) (ii) (iii) In view of Berge’s maximum theorem, the function is continuous in and there is a measurable mapping such that the maximum in (3.8) is achieved at [9, Theorem 2] or [10, Theorem 3.3]. Thus, the Markov policy is optimal for the horizon By induction, the Markov policy is optimal for the horizon This Markov policy is defined for Nature in the special MDP. It corresponds to the nonrandomized policy for Nature, which with a small abuse of notations we denote by the same letter
If the DM plays the deterministic risk-independent policy and Nature plays the nonrandomized policy then, according to Theorem 3.2,
@@
and for we have The last equality and (3.9) imply that the theorem is proved for
@@
where the first equality follows from unform -a.s. convergence the second inequality follows from the proved part of this theorem for and the last one follows from convergence of value iterations for the MDP with weakly continuous transition probabilities [9, Theorem 2] introduced in the proof of Theorem 3.2 when . ∎
The following corollary from Theorems 3.1 and 3.3 characterizes the optimal value of static CVaR in terms of the DRMDP.
@@
We recall that the DRMDP satisfies assumptions of an RMDP considered in Appendix A. Thus, the value functions exist and continuous in for finite and infinite time horizons , and uniformly in and There are nonrandomized optimal policies and persistently optimal policies for the DM and Nature. Theorem A.1 characterizes the sets of optimal policies and persistently optimal policies. In particular, for each player there exist optimal and persistently optimal nonrandomized Markov policies, and for each player has a deterministic optimal policy.

### v4

@@ Abstract
This paper studies optimization of Conditional Value-at-Risk (CVaR) for Markov Decision Processes (MDPs) with finite state and action sets. It introduces the Dynamically augmented CVaR (DCVaR) risk objective function and provides an algorithm for its optimization. This paper investigates a specially defined Robust MDP (RMDP), in which the state space is augmented with the tail risk level. This RMDP, which we call the Dynamically augmented RMDP (DRMDP), was introduced to the literature for calculations of optimal CVaR values by value iteration more than ten years ago, but, as was understood later, these value iterations compute lower bounds of minimal static CVaRs. DCVaR is defined as a time consistent version of the static CVaR, and it is a lower bound of the static CVaR. It also can be considered as a dynamic version of the nested CVaR. This paper provides an algorithm constructing a policy optimizing DCVaR of total discounted costs. The correctness of this algorithm is proved by studying a special mass transfer problem. The results on RMDPs needed for this paper are provided in the appendix.
Keywords: Markov decision process, conditional value at risk, optimal policy, algorithm
@@ 1 Introduction
Conditional Value-at-Risk (CVaR), also sometimes called Average Value-at-Risk (AVaR), is one of the most important and popular risk measures. For dynamic problems, it is possible to consider the so-called static CVaR, when the CVaR value is defined for each policy, and the goal is to find a policy optimizing this value. However, there are problems with this approach dealing with computational complexity of finding optimal policies and time inconsistency of static CVaRs. The issue of time consistency of risk measures is discussed in several works, including [22, 34, 35, 37, 38, 39]. An alternative approach is to consider the nested CVaR, which is defined via a Robust Markov Decision Pprocess (RMDP) whose state space coincides with the original state space of the problem. Another approach, which is also often used in practice, is to consider an RMDP, whose states are pairs consisting of the original state augmented with the risk level. In other words, the state is dynamically augmented by a number between 0 and 1 representing the tail risk level.
@@
The first group of the results in this paper establishes the relation between the static CVaR values and the DRMDP. First we observe that there is an optimal nonrandomized policy minimizing static CVaR, and, for a nonrandomized policy for a Markov Decision Process (MDP), its static CVaR is equal to the worst possible expected outcome if the Decision Maker (DM) plays this policy in the DRMDP. This observation clarifies the gap described in Hau et al. [16] and illustrates time inconsistency of the static CVaR for MDPs because, in order to implement this worst possible outcome, the second player in the DRMDP, whom we call Nature, may have to play a policy, which knows future decisions of the DM. A more natural assumption is that Nature plays its optimal policy, and this assumption leads to the definition of the DCVaR provided in this paper. This definition implies that value iterations proposed in Chow et al. [4] converge to minimal values of DCVaR.
@@
Classic methods for MDPs deal with risk-neutral objectives. In addition, there is a significant literature on risk-sensitive optimization initiated by Howard and Matheson [19] dealing mostly with optimization of expectations of utility functions rather than with optimization of expected total costs or average costs per unit time. Pioneering work by Markowitz [26] on mean-variance optimization for portfolio selection influenced research on MDPs with mean-variance criteria [11, 42], but variance of the total costs is not a convenient characteristic to deal with for sequential problems, and it is not a good risk measure, because variance is affected by losses and gains in the same way. CVaR is an attractive alternative to variance both because of its practical importance and rich mathematical properties; see [30, 31, 32, 40]. CVaR is a risk measure that has gained popularity in various applications including finance. It has the ability to safeguard a DM from the tail events by focusing on an average of the largest losses and by doing so it provides a more comprehensive view for risk management than threshold-based risk measures such as Value-at-Risk(VaR). CVaR is also known as expected shortfall in financial literature and has great importance in terms of financial regulation.
@@
We summarize the organization of this paper and its main results. Preliminary information on CVaR and MDPs is provided in Section 2. In Section 3 we prove the existence of a nonrandomized policy minimizing CVaR of the total discounted costs. Then we define the DRMDP as an RMDP in which the DM chooses actions, and Nature assigns tail risk levels. States of the DRMDP are states of the original MDP augmented with tail risk levels. Policies for the original MDP are particular policies of the DM in the DRMDP when the DM does not observe past and current tail risk levels, except for the initial one, and we call these policies risk-independent. We show that for a nonrandomized risk-independent policy played by the DM, CVaR of total discounted costs is equal to the worst expected total discounted costs the DM can incur by playing this policy in the DRMDP. This result illustrates time inconsistency of the static CVaR and implies that the minimal static CVaR value for the MDP is equal to the minimal value for the class of nonrandomized risk-independent policies in the DRMDP. This equality illustrates a positive gap between the optimal static CVaR value and the limit of value iterations. This gap was discovered in Hau et al. [16]. Section 4 defines a DCVaR, describes its properties, and introduces a DRMDP1, which is obtained from the DRMDP by modifying one-step costs and transition probabilities. The payoff and value functions for the DRMDP1 are equal to these functions for the DRMDP multiplied by tail risk levels. The sets of optimal policies for the DRMDP and DRMP1 coincide if the initial tail risk level is positive. The advantage of the DRMDP1 is that its value function is concave in the tail risk level.
@@
Section 5 introduces the algorithm for constructing a nonrandomized optimal policy minimizing DCVaR, which was announced in [5]. Sections 6–8 deal with proving that this algorithm constructs an optimal policy. Section 6 describes the properties of the mass transfer problem relevant to optimal solutions for Nature. Section 7 describes the properties of optimal values and properties of the sets of optimal actions. Section 8 proves correctness of the algorithm. Section 9 provides extensions to random cost. Appendix A provides results on RMDPs used in this paper.
@@ 2 Preliminaries: CVaR and MDPs
where is the distribution function of In addition, if this expectation exists, and where by definition; see [30, 31, 32] and [40, Chapter 6] for additional details regarding CVaR.
@@
where , and is the set of random variables on the probability space defined as and this set is called the dual CVaR risk envelope. In this paper, we sometimes write a constant instead of a function equal to this constant, and comparisons between random variables are usually understood -a.s.
Let be a sub -algebra of that is, is a -algebra on and A -measurable random variable is called a random tail risk level. Then the conditional CVaR given the -algebra and a random tail risk level is defined [29, Definition 17] as the -measurable random variable
where the essential supremum is taken with respect to the conditions probability and is the set of nonnegative -measurable random variables on such that and According to the CVaR decomposition theorem by Pflug and Pichler ([29], Lemma 22),
@@
An MDP is a model of a natural family of controlled stochastic sequences. It is a tuple where is a set of states, is a set of actions, and for each state there is a nonempty set of available actions is a real-valued cost function, and is a transition probability. The time is discrete, and, if at some time instance an action is selected at a state , then the system moves to the next state according to and the cost is collected, allowing the cost to depend on the next state. In this paper we always assume that and are finite sets.
@@
The risk-neutral approach is to consider the expected total discounted cost For a tail risk level and a finite horizon or an infinite horizon let us consider the objective function where , and is the random variable defined on the probability space in formulae (2.2) or (2.3). The value is sometimes called a static CVaR. Computing optimal policies minimizing the static CVaR is a hard problem [2, Remark 5.1].
Chow et al. [4] changed the original problem formulation with the state space to the formulation when the states of the problem consist of state-risk pairs where is the original state and is the tail risk level. In other words, the state space is augmented in [4] with the tail risk level, and the new state space is Motivated by the Pflug-Pichler CVaR decomposition theorem [29], Chow et al. [4] introduced the Robust MDP (RMDP) with the state space which is formally defined in Section 3 and called the Dynamically augmented RMDP (DRMDP). Infinite horizon was considered in [4], and a finite horizon version was considered in [23], where it was assumed that the DM knows tail risk levels at all steps. Hau et al. [16] showed that the value of DRMDP is a lower bound of the minimal CVaR values for the original MDP. In Section 4 we introduce the Dynamically augmented CVaR (DCVaR) objective function, which is a lower bound of CVaR, and the value of DRMDP is equal to the minimal value of DCVaR.
In the conclusion of this section, we recall two concepts for MDPs relevant to more general objective criteria than expected total costs: MDPs with general objective functions and abstract dynamic programming. According to [7], a general objective function is a function such that, if for and then This definition is consistent with the notion of a law-invariant risk measure [29, 40].
@@ 3 Static CVaR and Robust MDPs
The main results of this section are: (i) the existence of nonrandomized policies minimizing static CVaR (Theorem 3.1), and (ii) the minimal value of static CVaR is the minimal guaranteed expected total discounted payoff for the DM in the RMDP introduced in Chow et al. [4], if the DM plays policies from the MDP; see Corollary 3.4 below. This corollary provides a game-theoretic interpretation of the gap between the optimal value of CVaR and optimal value of this RMDP discovered in Hau et al. [16].
@@
be the optimal value of the static CVaR. The following theorem states the existence of nonrandomized optimal policies minimizing static CVaRs and convergence of finite-horizon optimal values of static CVaRs to the optimal infinite-horizon value. We recall that is fixed. We also recall that throughout the entire paper the discount factor if and if
@@
For every or for for the static CVaR optimization problem there exists a nonrandomized optimal policy for which for all In addition,
@@
We observe that it is sufficient to proof that for every there exists a nonrandomized optimal policy for the initial state Indeed, let be the set of all nonrandomized -horizon policies. If is optimal for an initial state then the nonrandomized policy is optimal, where Let us fix and prove the existence of Let us start with In this case the set is finite. Therefore, the strategic measure for an arbitrary policy can be presented as a convex combination of strategic measures for nonrandomized policies [6, Theorem 1]: there are numbers such that and Since taking CVaR of a mixture of distributions is a concave operation [28, Proposition 3.2],
@@
Let us consider the infinite-horizon problem. In the rest of this proof, let us set for all If we add a constant to the cost function then the all CVaR values for an -horizon problem will be changed by the constant Thus, this addition preserves sets of optimal policies. Since the cost function is bounded, without loss of generality we can assume that the function is nonnegative. Let us assume that and let is an upper bound of the cost function, that is, for all and
@@
Let us construct a nonrandomized optimal infinite-horizon policy Let be nonrandomized optimal -horizon policies. Since the sets and where are finite, it is possible to choose a sequence of integers such that for all and for all We define for all
Now let for some there is an increasing sequence of natural numbers such that for each history policies make the same decisions for all We select a subsequence of the sequence such that for each history all policies make the decision at the step which is the same for all By repeating this procedure and taking we construct a nonrandomized policy
@@
We recall that and denote by the -dimensional Euclidian space. Let us consider the RMDP introduced in Chow et al. [4], which we call the Dynamically augmented RMDP (DRMDP). The DRMDP is defined by a tuple where the state space is action space is uncertainty space is with being the number of states in action sets for the DM at states are uncertainty sets for Nature are
where, and, for
one-step costs for and transition probabilities for and where is the Dirac measure on the interval concentrated at the point A DRMDP is a particular case of the RMDP model described in Appendix A, when the sets set-valued mapping one-step cost function and transition probability have specific forms. In particular, states for DRMDPs are denoted by The sets of policies for the DM and Nature are denoted by and respectively. We shall use notation where and for the expected total discounted payoffs over the horizon with the terminal payoff which is continuous in As explained after formulae (A.2) and (A.4), the optimal values are continuous functions. Since is a finite set, this means that the functions are continuous in
We observe that that is, the set of all policies for the original MDP is the subset of the set of policies for the DM in the DRMDP. We call policies from risk-independent. The fundamental difference between a policy and a policy from is that the policy does not have information about the current and past tail risk levels except the initial tail risk level which we consider to be fixed.
The following theorem states that, for each risk-independent policy for the DM, Nature can play a policy maximizing DM’s losses. The proofs of Theorems 3.2 and A.6 are based on the well-known fact that, if in a two-player stochastic game a policy of one player is fixed, then another player deals with an MDP. In general, the states of this MDP are histories in the game. However, the state space of this MDP can be simplified, if the fixed policy belongs to a special class. For example, if the fixed policy is stationary, then another player deals with an MDP whose states are the states of the stochastic game. There are differences between the proofs of Theorems 3.2 and A.6. The proof of Theorem A.6 follows from the validity of sufficient conditions for the existence of optimal policies for MDPs with setwise continuous transition probabilities [36, 9], which takes place because all the sets are finite. The proof of Theorem 3.2 follows from the validity of sufficient conditions for the existence of optimal policies for MDPs with weakly continuous transition probabilities [36, 8], which takes place because the set and all the sets are finite and because of weak continuity of the transition probability The remaining statements of this section deal with the DRMDP.
@@
For each risk-independent policy for the DM and for each or there exists a nonrandomized policy which depends on and such that
@@
If the DM plays a risk-independent policy Nature deals with an MDP with states The set of available actions at each state is where and The one-step reward for Nature is and this function is bounded and continuous, where If an action is chosen at state then the next state is with the probability where the following conditions hold: (a) (b) (c) (d) Other transitions are impossible since The transition probability is weakly continuous in because is a bounded continuous function and and are distributions on finite sets depending on the finite numbers of conditions. We consider the expected total discounted rewards for this problem with the discount factor The initial state is where which is the initial tail risk level, and This MDP satisfies the weakly continuous conditions, which implies the existence of optimal Markov policies for finite-horizon problems and optimal deterministic policies for infinite-horizon problems [36] or [9, Theorem 2]. These optimal policies define the policy for Nature whose existence is stated in the theorem. ∎
@@
Let us consider the MDP defined at the beginning on the proof of Theorem 3.2 for We fix and compute an optimal policy for Nature. If at time Nature is at state of this MDP, then the optimal 1-step value for Nature is
where there is a nonrandomized optimal 1-step Markov policy [9, Theorem 2], and, in view of Berge’s maximum theorem, the function is continuous. This means that this function is continuous in since all other variables in take a finite number of values. Nature chooses the optimal action without using any knowledge of the previous states and actions. This is true for because the optimal value of does not depend on the distributions of future actions and states following the terminal state
Let us make the induction assumption that for some integer there is a Markov policy such that this policy is optimal for the horizon and the value function is continuous, which means that it is continuous in Then
for where and In view of Berge’s maximum theorem, the function is continuous in and there is a measurable mapping such that the maximum in (3.4) is achieved at [9, Theorem 2] or [10, Theorem 3.3]. Thus, the Markov policy is optimal for the horizon By induction, the Markov policy is optimal for the horizon This Markov policy is defined for Nature in the special MDP. It corresponds to the nonrandomized policy for Nature, which with a small abuse of notations we denote by
If the DM plays the nonrandomized risk-independent policy and Nature plays the nonrandomized policy then, according to Theorem 3.2,
@@
and, for and we have The last equality and (3.5) imply that the theorem is proved for
@@
where the first equality follows from unform -a.s. convergence the second equality follows from the proved part of this theorem for and the last one follows from convergence of value iterations for the MDP with weakly continuous transition probabilities [9, Theorem 2] introduced in the proof of Theorem 3.2 when ∎
The following corollary from Theorems 3.1 and 3.3 characterizes the optimal value of static CVaR in terms of the DRMDP. Recall that is the set of policies for the original MDP.
@@
We recall that the DRMDP satisfies assumptions of an RMDP considered in Appendix A. Thus, the value functions exist and are continuous in for finite and infinite time horizons , and as uniformly in and Theorem A.4 implies the existence of optimal policies and characterizes some classes of optimal policies. In particular, for each player there exist optimal nonrandomized Markov policies, and for each player has a deterministic optimal policy.
Let us consider the gap between the optimal value of CVaR and the value of the RMDP. For or for and for
where the second equality follows from Corollary 3.4 and Theorem A.7, and the inequality follows from It is shown in [16] that this gap can be positive for see also [13].
If then we deal with a one-step problem, which can be solved easily by performing one step of value iterations for the DRMDP with In this case, the gap is equal to 0. Therefore, in the rest of this paper, except Appendix A, we consider